Kaymarie08:

Richard can read 1/4 of a book in 4/5 of an hour. At this rate, how much can Richard read in one hour

Convert:

1/4 = 4/5 hours (1/4) 5 = (4/5) 5 5/4 = 20/5 5/4 = 4 hours 5/4 divided by 4 = 4 hours divided by 4 5/4 multiplied by 1/4 = 4 multiplied by 1/4 0.3125 = 1 hour 5/16 = 1 hour Answer: Richard can read 5/16 of the book in 1 hour.